### v1

# Self-organizing Multiagent Target Enclosing under Limited Information and Safety Guarantees

Abstract
This paper introduces an approach to address the target enclosing problem using non-holonomic multiagent systems, where agents autonomously self-organize themselves in the desired formation around a fixed target. Our approach combines global enclosing behavior and local collision avoidance mechanisms by devising a novel potential function and sliding manifold. In our approach, agents independently move toward the desired enclosing geometry when apart and activate the collision avoidance mechanism when a collision is imminent, thereby guaranteeing inter-agent safety. We rigorously show that an agent does not need to ensure safety with every other agent and put forth a concept of the nearest colliding agent (for any arbitrary agent) with whom ensuring safety is sufficient to avoid collisions in the entire swarm. The proposed control eliminates the need for a fixed or pre-established agent arrangement around the target and requires only relative information between an agent and the target. This makes our design particularly appealing for scenarios with limited global information, hence significantly reducing communication requirements. We finally present simulation results to vindicate the efficacy of the proposed method.
Index Terms:
Multiagent Systems, Target Enclosing, Collision Avoidance, Motion Planning, Safety, Self-organizing Behavior.
I Introduction
In recent times, autonomous vehicles have gained significant prominence in a range of tasks, including reconnaissance, surveillance, crop and forest monitoring, and search and rescue missions, e.g., [1, 2, 3, 4, 5]. Most of these tasks share a common behavior where an autonomous vehicle, known as the pursuer, monitors an object of interest, known as the target. In general, the pursuer’s stable motion around the target by maintaining the desired separation from the target is referred to as target enclosing [6] or encirclement/circumnavigation/entrapment when the enclosing orbit is a circle [7].
One of the earliest approaches for target circumnavigation entails pursuers confining the target in a specific geometric shape posing as a formation control problem. The authors in [8, 9] utilize a cyclic pursuit strategy to cooperatively enclose a target, achieving and maintaining uniform angular spacing among the agents. Vector field guidance is another popular approach, where vector fields can be designed to create a limit cycle or periodic orbit around the target, guiding agents to desired geometric shapes [10, 11]. However, each orbit may require a separate field, thus limiting its applicability in a multiagent scenario. Several other guidance strategies have also been developed for trapping targets in circular orbits, based on differential geometry [12], model predictive control [13], and reinforcement learning [14]. Many of the works in the target enclosing have focused on reducing the information required for developing the control laws. For example, some strategies utilize only range-only measurements [15, 16, 17], bearing angle measurement [18, 19], range and bearing information [20, 7, 21]. The general moving target enclosing problem using a single agent was studied in [22, 23], whereas the authors in [24] guided multiple agents to the desired proximity with equal angular spacing. Note that these works did not account for explicit inter-agent safety.
Most studies on target enclosing have focused on ensuring that pursuers reach the target without considering their safety or coordinating their movements with other pursuers. However, this approach can result in collisions between pursuers, which may result in mission failure. To address this issue, some studies have introduced decentralized methods for collision avoidance between pursuers, allowing them to orbit the target while maintaining a safe distance from each other [8, 9, 13, 24, 25, 26]. However, these methods often rely on strict formations, such as equal angular spacing, which can limit the pursuers’ movement and increase the complexity of the guidance strategy. By eliminating the need for rigid formations and reducing the number of agents that each pursuer must communicate with, we can simplify the guidance strategy and improve its robustness to changes in the number of pursuers in the enclosing formation.
Motivated by the aforementioned studies, our work aims to create efficient target-enclosing guidance laws that do not require extensive information exchange or predefined agent configurations. Our main contributions include developing a distributed guidance law for pursuers to safely enclose a target while avoiding collisions, using a minimal amount of information, and no fixed agent formations. Each pursuer only coordinates with one other pursuer when necessary to prevent collisions, thereby ensuring inter-agent safety. Our approach introduces a novel decision-making protocol for selecting the appropriate pursuer for coordination and designing a guidance strategy to ensure collision avoidance. We design the guidance law based on a custom potential function that combines attractive and repulsive forces to guide the pursuers toward the desired orbit around the target safely and efficiently. We also demonstrate that our proposed guidance law is robust to the motion of other pursuers and requires only relative information, potentially eliminating the need for costly agent-to-agent communication.
II Problem Formulation
II-A Kinematics of Agents’ Relative Motions
Consider a multiagent system consisting of $n$ pursuers and a single stationary target $T$. We use $\mathcal{L}$ to denote the set of the pursuer, that is, =$\mathcal{L}=\{P_{1},P_{2},\ldots,P_{n}\}$. Figure 1 depicts the engagement geometry between the pursuers and the target, where $[x_{i},y_{i}]^{\top}\in\mathbb{R}^{2}$ denotes the position of the $i$th pursuer in the inertial frame of reference, and $\chi_{i}\in[-\pi,\pi)$ denotes its heading angle. Assuming the vehicles to be point masses, the motion of the $i$th pursuer is governed by
$\dot{x}_{i}=v\cos\chi_{i},\quad\dot{y}_{i}=v\sin\chi_{i},\quad\dot{\chi}_{i}=\frac{a_{i}}{v},$
where $v\in\mathbb{R}_{>0}$ denotes the common and constant speed for all pursuers and $a_{i}$ denotes the lateral acceleration, also the control input, of the $i$th pursuer. This model, albeit simple, is practical when dealing with vehicles that are turn-constrained and utilize lateral forces (e.g., lift and side force) to maneuver, such as fixed-wing aircraft, missiles, and underwater vehicles [27, 6, 22, 7, 21, 1]. Here we assume that $|a_{i}|\leq a_{i}^{\max}$.
Figure 1: Target enclosing scenario in the inertial frame of reference.
The engagement geometry between the vehicles in the relative frame of reference is depicted in Figure 2, where $r_{ij}$ denotes the distance between the $i$th and the $j$th pursuer ($i\neq j$) and accordingly $r_{iT}$ signifies the distance with the target. Similarly, $\theta_{iT}$ denotes the line-of-sight (LOS) angle for the $i$th pursuer with respect to the target and $\theta_{ij}$ is the LOS angle between the $i$th and the $j$th pursuer referenced from the $i$th pursuer. The bearing angle of the $i$th pursuer with respect to the target is $\sigma_{i}=\chi_{i}-\theta_{iT}$. Thus, the kinematics of the relative motion for the multiagent system is given by
$\dot{r}_{iT}$ $=-v\cos\sigma_{i}$ (1a)
$\dot{\theta}_{iT}$ $=-\frac{v\sin\sigma_{i}}{r_{iT}}.$ (1b)
These equations are obtained by projecting the $i$th pursuer’s relative velocities in the directions parallel/normal to the $i$th pursuer-target LOS.
Figure 2: Target enclosing scenario in the relative frame of reference.
II-B Direction/Sense of Target Enclosing
For the target-centric relative system in (II-A), $r_{iT}$ is measured from $P_{i}$ to target and $\theta_{iT}$ is measured in a counter-clockwise sense from the horizontal reference. Therefore, positive angular rates pertain to the counter-clockwise revolution of the pursuers about the target. Further, it is evident from (1b) that $\dot{\theta}_{iT}<0,\;\forall\;\sigma_{i}\in(0,\pi)$, $\dot{\theta}_{iT}>0,\forall\,\sigma_{i}\in(-\pi,0)$ and $\dot{\theta}_{iT}=0,\forall\,\sigma_{i}\in\{0,\pm\pi\}$.
Remark 1.
One may notice from these observations that the $i$th pursuer revolves around the target in a counter-clockwise direction when $\dot{\theta}_{iT}>0$ and vice-versa. From the bearing angle relationship, it follows that the positive values of $\sigma_{i}$ represent the scenarios when the $i$th pursuer’s heading angle leads the LOS angle in the counter-clockwise direction. Therefore, if $\sigma_{i}$ is assumed to be constrained within $(0,\pi)$, then $\dot{\theta}_{iT}<0$ can be ensured for all $t\geq 0$. This aspect is critical from a safety perspective because it limits the $i$th pursuer’s revolution to the clockwise direction only around the target. This helps ensure that the pursuers do not go in opposite senses while converging on a circle of the same radius, which could lead to a collision.
To describe the relationship between the $i$th and the $j$th pursuer at any instant of time in terms of their relative positions, let us define the terms inter-agent separation $d_{ij}$ and the inter-agent angular spacing $\psi_{ij}$ as
$d_{ij}$ $=r_{iT}-r_{jT},$ (2a)
$\psi_{ij}$ $=\theta_{iT}-\theta_{jT},$ (2b)
where $\psi_{ij}\in[-\pi,\pi)$, as depicted in Figure 3, where the agents (the pursuers) are enclosing the target in a clockwise direction. In Figure 3, the dotted circles represent the loiter circles around the target with a radius equal to the instantaneous pursuer-target distance. The positive values of $d_{ij}$ imply that the $i$th pursuer is on a larger loiter circle compared to the $j$th pursuer. Hence, $d_{ij}>0$ implies $\dfrac{{r}_{iT}}{r_{jT}}>1$, and $d_{ij}<0$ implies that $10<\dfrac{{r}_{iT}}{r_{jT}}<$ 1. Similarly, the positive values of $\psi_{ij}$ result in a scenario when $\theta_{jT}$ leads $\theta_{iT}$ in a clockwise sense. The red arrows in Figure 3 denote the positive conventions of $d_{ij}$ and $\psi_{ij}$ considered in this paper. It is also worth noting that for any two agents $P_{i}$ and $P_{j}$, $d_{ij}$ and $\psi_{ij}$ together represent the inter-agent radial distance $r_{ij}$. From Figure 2, using the cosine angle formula from triangle $P_{i}-P_{j}-T$, we can obtain the distance between $P_{i}$ and $P_{j}$ as, $r_{ij}^{2}=r_{iT}^{2}+r_{jT}^{2}-2r_{iT}r_{jT}\cos\psi_{ij}$, which reduces to,
$r_{ij}^{2}=2r_{jT}^{2}\left(1-\cos\psi_{ij}\right)+d_{ij}^{2}+2r_{jT}d_{ij}\left(1-\cos\psi_{ij}\right),$ (3)
on using $r_{iT}=r_{jT}+d_{ij}$ from (2a). Hence, if both $d_{ij}\to 0$ (coincident loiter circles) and $\psi_{ij}\to 0$ (coincident pursuer-target LOS), then it essentially implies that $r_{ij}\to 0$, which means inter-agent collision since more than one pursuer occupies the same position.
Figure 3: Inter-agent reference geometry.
II-C Design of Control Objectives
We aim to make the pursuers organize themselves in a circular orbit around the target while also ensuring that they do not collide with each other, as the orbit is the same for all of them. To solve the multiagent target enclosing problem under limited information and safety constraints, we need to ensure that for each pursuer, starting from any feasible initial conditions, the following holds:
$\lim_{t\to\infty}\left|r_{iT}-r_{d}\right|\to 0,$ (4)
where $r_{d}$ denotes the fixed desired proximity of the pursuers from the target, which is the same for every pursuer. It is worth noting that fulfilling the objective in (4) is not sufficient to guarantee inter-agent safety because the circular orbit is the same for all pursuers during target enclosing. In addition to (4), the requirement that $r_{ij}>0,\,\forall\,i\neq j,$ needs to be satisfied throughout the engagement for safety. This condition ensures that the inter-agent spacing is always greater than a positive value, that is, no two pursuers converge to the same point on the desired enclosing orbit.
Remark 2.
As the number of agents increases, exhaustively checking the condition of inter-agent spacing for safety ($r_{ij}>0$) leads to computational and design complexities in terms of sensing and communication. For an $i$th agent, there exists some $r_{ij}$ such that checking their sign to ensure safety would be tedious and redundant. That said, an agent may require extensive sensing and communication with every other agent (not just its neighbors), which may not be efficient.
Figure 4: Engagement between pursuers without safety considerations.
To elucidate the implications of Remark 2, let us analyze a typical case of target enclosing involving two pursuers, as shown in Figure 4. If $P_{j}$ is initially located closer to the target than $P_{i}$ such that $d_{ij}>0$ and $\psi_{ij}<0$, then $\dot{\theta}_{iT}<\dot{\theta}_{jT}$. This essentially means that the component of the relative velocity of $P_{j}$ normal to $r_{jT}$ is greater than that of $P_{i}$, thereby making $\dot{\psi}_{ij}>0$. This leads to coincident pursuer-target LOS between $P_{i}$ and $P_{j}$ at some instant of time during target enclosing. Similarly, the component of the relative velocity of $P_{j}$ along $r_{jT}$ will be less than that of $P_{i}$, thereby making $\dot{d}_{ij}<0$. This results in the pursuers being on the same loiter circle at some instant of time during target enclosing. Consequently, a collision between $P_{i}$ and $P_{j}$ may eventually occur, as demonstrated in Figure 4, where ${(\cdot)}^{{}^{\prime}}$ and ${(\cdot)}^{{}^{\prime\prime}}$ represent the variables related to the pursuers as time progresses since both $d_{ij}\to 0$ and $\psi_{ij}\to 0$ as a result of $\dot{d}_{ij}<0$ and $\dot{\psi}_{ij}>0$. Therefore, one may speculate that inter-agent collisions can be prevented if such conditions/scenarios never arise during target enclosing.
The previous analysis showed that collisions between pursuers occur only under specific circumstances. Our method takes advantage of this fact to restrict each pursuer’s motion to avoid collisions solely with those pursuers that have the potential to collide in the future. It is important to note that when there are multiple pursuers, several of them may collide with $P_{i}$. Consequently, we devise strategies to prevent collisions with a single pursuer while taking into account the potential collisions with all the other pursuers. To this end, we first identify the set of all pursuers that are within the sensing range of the $i$th pursuer (i.e., $r_{ij}\leq r_{s}$) and may potentially collide with it. Let us refer to this set as
$\mathcal{N}_{i}\coloneqq\left\{P_{j}\in\mathcal{L}\,\big{|}\ d_{ij}>0\,\&\,\psi_{ij}<0\right\},~{}\forall\,i\neq j,$ (5)
which represents the set of all colliding neighbors of $P_{i}$. Out of all such pursuers in $\mathcal{N}_{i}$, we identify those pursuers that are situated on a loiter circle whose radius is closest (just smaller) to the one on which $P_{i}$ is located. Let’s define
$\mathcal{Z}_{i}\coloneqq\left\{P_{j}\in\mathcal{N}_{i}\big{|}\,d_{ij}=\min\limits_{P_{k}\in\mathcal{N}_{i}}d_{ik}\right\},$ (6)
for any $k$th pursuer such that $k\neq j$. Note that the sets $\mathcal{N}_{i}$ and $\mathcal{Z}_{i}$ may contain more than one pursuer. This happens when multiple neighboring pursuers belonging to the set $\mathcal{Z}_{i}$ share the same distance from the target. We, however, aim to develop a self-organizing guidance strategy under limited information and safety constraints. This would entail limiting the communication among the pursuers. Toward this objective, we define another set
$\mathcal{C}_{i}\coloneqq\left\{P_{j}\in\mathcal{Z}_{i}\big{|}\,\psi_{ij}=\max\limits_{P_{k}\in\mathcal{N}_{i}}\psi_{ik}\right\},$ (7)
that selects only one potentially colliding neighbor from the set $\mathcal{Z}_{i}$, thereby allowing us to analyze the behavior of only one pair of pursuers at any instant of time. We refer to the set $\mathcal{C}_{i}$ as the set of the nearest colliding pursuer. In essence, $\mathcal{C}_{i}\subset\mathcal{Z}_{i}\subset\mathcal{N}_{i}$.
Remark 3.
Based on the conditions given in (6) and (7), the set $\mathcal{C}_{i}$ contains at most one neighboring pursuer since a unique combination of $(d_{ij},\psi_{ij})$ represents a distinct position of $P_{j}$ (the nearest colliding pursuer) about $P_{i}$. Therefore, it is sufficient to prevent inter-agent collision between any such pair, thereby ensuring safety under limited information.
In this work, we show that if $d_{ij}>0$ for all $t\geq 0$ in addition to (4), then the pursuers will exhibit a self-organizing behavior during target enclosing while also guaranteeing safety. Ensuring $d_{ij}>0$ will imply that $r_{ij}>0$ even if $\psi_{ij}\to 0$, meaning that every pursuer will maintain a safe distance with respect to every other pursuer in its vicinity.
The scenario represented in Figure 5 considers three pursuers, $P_{i}$, $P_{j}$, and $P_{k}$ aiming to enclose the target within a circle while also ensuring safety. The dotted curves in Figure 5 represent their instantaneous loiter circles around the target. In this scenario, $P_{j}$ and $P_{k}$ initially avoid encountering a nearest colliding pursuer, thus converging to the desired proximity in the same order as they started. Meanwhile, $P_{i}$ encounters $P_{j}$ as its nearest colliding pursuer, prompting $P_{i}$ to maintain $d_{ij}>0$ to avoid collision with $P_{j}$. Subsequently, at a later time during enclosing, $\psi_{ij}\to 0$, enabling $P_{i}$ to relax its safety constraint and settle behind $P_{j}$, ultimately resulting in collision-free self-organizing target enclosing.
Figure 5: Multagent target enclosing scenario with safety constraints.
This observation facilitates us to formulate the safety constraint for the $i$th pursuer as the one of ensuring
$d_{ij}>0.$ (8)
This condition guarantees that all pursuers move along a loiter circle with a radius greater than its distance from the nearest colliding pursuer around the target. As a result, $\psi_{ij}\to 0$, but $d_{ij}\not\to 0$, which implies that $r_{ij}\not\to 0$, ensuring that each pursuer keeps a safe distance from its nearest colliding pursuer and is influenced by at most one other pursuer during enclosing.
Figure 5 shows the behavior of pursuers who ensure safety conditions in (8), where $P_{j}$ is the nearest colliding pursuer for $P_{i}$. Initially, $\psi_{ij}$ starts at a negative value and becomes positive over time, which is similar to the scenario without safety conditions presented in Figure 4. However, due to the safety condition, $P_{i}$ maintains a larger distance from the target than $P_{j}$, resulting in a positive $d_{ij}$. Once $\psi_{ij}=0$, $P_{j}$ is no longer the nearest colliding pursuer for $P_{i}$, and $P_{i}$ begins to converge to the desired enclosing geometry while $\psi_{ij}$ becomes increasingly positive. At steady state, $P_{i}$ autonomously converges to a finite positive value of $\psi_{ij}$ with respect to $P_{j}$. This behavior, where pursuers converge to a circle of desired proximity from the target while maintaining a safe distance from each other, is referred to as self-organizing multiagent target enclosing behavior in our paper. Unlike previous works [8, 9, 13, 24, 25, 26], the safety condition in (8) constrains $d_{ij}$ rather than $\psi_{ij}$, allowing each pursuer to independently converge to any finite $\psi_{ij}$ with respect to its neighbors. Such behavior enables pursuers to operate without the need for a predetermined formation structure, reducing the amount of information required and coordinating efforts with just one other pursuer, which is the nearest colliding pursuer. This affords pursuers greater autonomy to maneuver independently while still maintaining a stable formation/enclosing geometry. Moreover, pursuers can enter or exit the enclosing formation seamlessly, without disturbing the overall structure.
We are now in a position to summarize the control objectives in regards to the self-organizing multiagent target enclosing under limited information and safety guarantees, that is, $\lim_{t\to\infty}\left|r_{iT}-r_{d}\right|\to 0$, and $d_{ij}>0,~{}\forall\,P_{j}\in\mathcal{C}_{i}$.
III Self-organizing Multiagent Target Enclosing Guidance Law with Inherent Safety
In this section, we examine the interactions between pursuers and derive the criteria for identifying the nearest colliding pursuer. We then show that satisfaction of the safety condition in (8) is sufficient for developing a guidance strategy that steers all pursuers on the desired enclosing geometry around the target while also avoiding collisions.
III-A Analysis of Inter-Pursuer Behavior
Note that the LOS of the $i$th pursuer with respect to the target and the line that is normal to it at the $i$th pursuer’s position (the tangent line to its instantaneous loiter circle) partitions the engagement zone into four distinct regions that can be completely characterized using $d_{ij}$ and $\psi_{ij}$, as shown in Figure 6. Based on this partition, we can analyze the safety and behaviors of the pursuers as they maneuver to enclose the target. These four distinct region can be characterized as,
Region I: $d_{ij}>0,\psi_{ij}>0,~{}~{}\text{Region II:}\;$: $d_{ij}\leq 0,\psi_{ij}>0,$
Region III: $d_{ij}\leq 0,\psi_{ij}<0,~{}~{}\text{Region IV:}\;$: $d_{ij}>0,\psi_{ij}<0.$ (9)
Figure 6: Regions around the $i$th pursuer.
Lemma 1.
Pursuers situated in Regions I and III corresponding to the $i$th pursuer’s position do not have a tendency to collide with the $i$th pursuer.
Proof.
To demonstrate this, we examine the motion of pursuers in the four regions defined in (III-A) and determine the conditions under which both $d_{ij}\to 0$ and $\psi_{ij}\to 0$ can occur. We do this by deriving the dynamics of the inter-pursuer radial distance and inter-pursuer angular spacing, which are given by,
$\dot{d}_{ij}$ $=-v\left(\cos\sigma_{i}-\cos\sigma_{j}\right),$ (10a)
$\dot{\psi}_{ij}$ $=-v\left(\frac{\sin\sigma_{i}}{r_{iT}}-\frac{\sin\sigma_{j}}{r_{jT}}\right),$ (10b)
obtained by differentiating (II-B) with respect to time and using (II-A). Provided that all pursuers initiate their movement away from the target and tend to move toward it to enclose the latter, we can deduce from (1a) that $\dot{r}_{iT}<0,\dot{r}_{jT}<0$. This pertains to the scenario when they are moving toward the target, thus limiting their look angles within the interval $(0,\pi/2)$. Now, if a collision is imminent, then $d_{ij}\to 0$ and $\psi_{ij}\to 0$. Let us analyze each of these conditions in the four regions to analyze the situation when every pursuer is heading toward the target to enclose it.
Using (10a), we can determine the conditions that lead to $d_{ij}\to 0$ in the four regions as follows:
Region I: $\dot{d}_{ij}<0,\text{if}\,\sigma_{i}<\sigma_{j},$
Region II: $\dot{d}_{ij}>0,\text{if}\,\sigma_{i}>\sigma_{j},$
Region III: $\dot{d}_{ij}>0,\text{if}\,\sigma_{i}>\sigma_{j},$
Region IV: $\dot{d}_{ij}<0,\text{if}\,\sigma_{i}<\sigma_{j}.$ (11)
Next, we analyze cases when ${\psi}_{ij}\to 0$ when the above conditions for $d_{ij}\to 0$ also hold. We denote $\sigma_{i}^{\star}$ as the look angle about which ${\psi}_{ij}$ changes its sign, given by
$\sin\sigma_{i}^{*}=\frac{r_{iT}}{r_{jT}}\sin\sigma_{j},$ (12)
obtained by equating $\dot{\psi}_{ij}$ to zero in (10b).
In Regions I and IV, we have $d_{ij}>0$ or $\dfrac{r_{iT}}{r_{jT}}>1$, which implies $\sigma_{i}^{*}>\sigma_{j}$. Thus, $\sigma_{i}<\sigma_{j}$ leads to $\sin\sigma_{i}<\dfrac{r_{iT}}{r_{jT}}\sin\sigma_{j}$, resulting in $\dot{\psi}_{ij}>0$. Similarly, in Regions II and III, we have $\sigma_{i}^{*}<\sigma_{j}$, and hence $\sigma_{i}>\sigma_{j}$ implies $\sin\sigma_{i}>\dfrac{r_{iT}}{r_{jT}}\sin\sigma_{j}$, resulting in $\dot{\psi}_{ij}<0$. Therefore, we can summarize the inferences from the current analysis and those in (III-A) as
Region I: $\dot{d}_{ij}<0,\dot{\psi}_{ij}>0,\text{if}\,\sigma_{i}<\sigma_{j},$
Region II: $\dot{d}_{ij}>0,\dot{\psi}_{ij}<0,\text{if}\,\sigma_{i}>\sigma_{j},$
Region III: $\dot{d}_{ij}>0,\dot{\psi}_{ij}<0,\text{if}\,\sigma_{i}>\sigma_{j},$
Region IV: $\dot{d}_{ij}<0,\dot{\psi}_{ij}>0,\text{if}\,\sigma_{i}<\sigma_{j}.$ (13)
It follows from (III-A) and (III-A) that inter-pursuer collision can only occur in Regions II and IV during target enclosing. Therefore, no agents in Regions I and III will collide with $P_{i}$. This concludes the proof. ∎
Lemma 1 shows that the nearby pursuers in Regions I and III will not collide with $P_{i}$, so ensuring safety with those pursuers is unnecessary for $P_{i}$. This could reduce computational complexity and make the design simpler. Now, we explain why choosing one pursuer from the surrounding pursuers according to (7) is necessary to ensure safety among all pursuers.
Theorem 1 (Inter-pursuer Safety).
The selection of a single nearest colliding pursuer according to (7) and fulfilling the safety condition in (8) is sufficient to guarantee collision avoidance during target enclosing.
Proof.
Figure 7: Engagement with neighboring pursuers in Region II of $P_{i}$.
The essence of this theorem lies in the fact that for any two pursuers, if $d_{ij}>0$, then the $i$th pursuer only needs to ensure strict inter-pursuer safety with its nearest colliding neighbor, which can only be present in its Region II. According to the results in Lemma 1, $P_{i}$ can only collide with the pursuers located in Regions II and IV, and thus the pursuers located in these regions have a tendency to collide with $P_{i}$ if safety is not ensured. We show that identifying $P_{i}$’s nearest colliding neighbor according to (7) enables it to selectively choose a single pursuer from a larger group situated in Region IV, thereby enforcing safety guarantees in the self-organizing swarm, that is, $r_{ij}>0$ for $i\neq j$.
Let us consider a situation where pursuer $P_{i}$ has another pursuer, $P_{j}$, located within its Region II (denoted as Region Ii), as shown in Figure 7. Since $P_{i}$ is the nearest colliding pursuer for $P_{j}$, according to (7), $P_{j}$ must meet the safety criterion in (8) and maneuver accordingly. Hence, $P_{i}$ can safely approach the desired proximity without colliding with pursuers in its Region IIi. This means that $P_{i}$ can focus on avoiding collisions with pursuers in its Region IVi.
Before considering the second scenario, we first analyze the inter-pursuer distance $r_{ij}$ between $P_{i}$ and $P_{j}$ to arrive at the nearest colliding pursuer selection criterion discussed in (III-A). Following (III-A), we find the values of $d_{ij}$ and $\psi_{ij}$ that minimizes $r_{ij}$ by evaluating the partial derivatives of $r_{ij}^{2}$ as in (3) with respect to $d_{ij}$ and $\psi_{ij}$. We obtain the gradient $\nabla r_{ij}^{2}$ and Hessian $\nabla^{2}r_{ij}^{2}$ as,
$\nabla r_{ij}^{2}=$ $\begin{bmatrix}\dfrac{\partial r_{ij}^{2}}{\partial d_{ij}}\\ \dfrac{\partial r_{ij}^{2}}{\partial\psi_{ij}}\end{bmatrix}=\begin{bmatrix}2d_{ij}+2r_{iT}\left(1-\cos\psi_{ij}\right)\\ 2r_{iT}^{2}\sin\psi_{ij}+2r_{iT}d_{ij}\sin\psi_{ij}\end{bmatrix},$ (18)
$\nabla^{2}r_{ij}^{2}=$ $\begin{bmatrix}2&2r_{iT}\sin\psi_{ij}\\ 2r_{iT}\sin\psi_{ij}&2r_{iT}^{2}\cos\psi_{ij}+2r_{iT}d_{ij}\cos\psi_{ij}\end{bmatrix},$ (21)
It can be readily observed from (3), (18) and (21) that the minimum value of $r_{ij}=0$ is attained only when $d_{ij}=0$ and $\psi_{ij}=0$, since at these values $\nabla r_{ij}^{2}=0$ and $\nabla^{2}r_{ij}^{2}\succ 0$ (positive definite).These inferences follow the discussions in Section II-B. From (III-A), we know that in Region IV of $P_{i}$ (Region IVi in Figure 7), $d_{ik}>0$ and $\psi_{ik}<0$ for an arbitrary pursuer $P_{k}$ located therein. Therefore, when $r_{ik}\neq 0$, choosing $P_{k}$ with the smallest $d_{ik}$ and largest $\psi_{ik}$ will select only one pursuer that is closest to $P_{i}$ in terms of inter-pursuer distance. Our approach ensures safety among pursuers and allows them to self-organize during target enclosing by first picking pursuers located on the nearest loiter circle around the target to minimize $d_{ij}$. This set of pursuers is denoted by $\mathcal{Z}_{i}$, as seen in (6). Next, we select the nearest colliding pursuer with the largest $\psi_{ij}$ from the set $\mathcal{Z}_{i}$, fulfilling the conditions in (7). A pseudocode representation of the algorithm for each pursuer to identify the nearest colliding pursuer is provided in Algorithm 1.
Algorithm 1 Select Nearest Colliding Pursuer
1:Input: Set of pursuers $\mathcal{L}$, sensing range $r_{s}$, pursuer $P_{i}$
2:Output: Set of nearest colliding pursuer $\mathcal{C}_{i}$
3:Initialize $\mathcal{N}_{i}\leftarrow\{\}$, $\mathcal{Z}_{i}\leftarrow\{\}$, $\mathcal{C}_{i}\leftarrow\{\}$
4:for each pursuer $P_{j}\in\mathcal{L}$ do
5:     if $r_{ij}\leq r_{s}$ and $d_{ij}>0$ and $\psi_{ij}<0$ then
6:         Add $P_{j}$ to $\mathcal{N}_{i}$
7:     end if
8:end for
9:for each pursuer $P_{j}\in\mathcal{N}_{i}$ do
10:     if $d_{ij}=\min_{P_{k}\in\mathcal{N}_{i}}d_{ik}$ then
11:         Add $P_{j}$ to $\mathcal{Z}_{i}$
12:     end if
13:end for
14:Select $P_{j}\in\mathcal{Z}_{i}$ such that $\psi_{ij}=\max_{P_{k}\in\mathcal{N}_{i}}\psi_{ik}$
15:Add $P_{j}$ to $\mathcal{C}_{i}$
In the second situation, we consider the presence of multiple pursuers ($P_{j}$, $P_{k}$ and $P_{l}$) in Region IVi, as shown in Figure 8. Given the scenario, pursuer $P_{j}$ has the shortest distance to $P_{i}$ ($r_{ij}$), while pursuer $P_{k}$ has the smallest distance between itself and another pursuer ($d_{ik}$) based on the loiter circles. Additionally, pursuer $P_{l}$ has the largest inter-pursuer angular spacing between itself and $P_{i}$ ($\psi_{il}$). This implies $d_{ik}<d_{ij}<d_{il}$, and $\psi_{il}>\psi_{ij}>\psi_{ik}$. From (7), we can infer that $P_{k}$ is the nearest colliding pursuer for $P_{i}$, who only needs to concern itself with avoiding collision with $P_{k}$ by executing a suitable maneuver.
For the sake of contradiction, let us consider either $P_{j}$ or $P_{l}$ to the nearest colliding pursuer for $P_{i}$. First, considering $P_{j}$ to be the nearest colliding pursuer for $P_{i}$, we have $d_{ij}>0$ based on condition (8), which also implies $d_{il}>0$. However, this will not ensure $d_{ik}>0$ and result in a collision between $P_{i}$ and $P_{k}$. Similarly, we can show that selecting $P_{l}$ as the nearest colliding pursuer will not ensure the prevention of collision of $P_{i}$ with $P_{j}$ and $P_{k}$, since $d_{il}>0$ does not imply $d_{ij}>0$ and $d_{ik}>0$. Therefore, $P_{k}$ is the appropriate choice of nearest colliding pursuer chosen as per condition (7), which requires less information than ensuring safety with a pursuer other than the nearest colliding one (which will require additional conditions). By guaranteeing safety condition (8) with respect to $P_{k}$, we can infer that $d_{ik}>0$. This subsequently implies that $d_{ij}>0$ and $d_{il}>0$, since $P_{k}$ is closer to $P_{i}$ than $P_{j}$ and $P_{l}$. In other words, ensuring that $P_{i}$ avoids colliding with $P_{k}$ automatically ensures that $P_{i}$ avoids colliding with $P_{j}$ and $P_{l}$ as well. This concludes the proof.
Figure 8: Engagement with neighboring pursuers in Region IV of $P_{i}$.
Every pursuer starting away from the target may encounter a nearest colliding pursuer before reaching the desired proximity around the target. Under such a case, the $i$th pursuer should maneuver to only avoid collision with its nearest colliding pursuer, ensuring safety conditions in (8). Consequently, this allows the pursuers to settle behind the nearest colliding pursuer on the enclosing orbit, allowing the pursuers to self-organize around the target with inherent safety.
Remark 4.
Our methodology capitalizes on the inherent self-organization behavior of each pursuer to adjust their individual trajectory without necessitating explicit cooperation or prior coordination among them. This decentralized approach allows pursuers to operate independently, adapting to changing environmental conditions and avoiding potential collisions while still working towards the collective goal of enclosing the target. By eliminating the need for rigid formations or vehicle-to-vehicle communication, our solution offers a more practical and robust alternative to traditional centralized control strategies.
III-B Design of Potential Function
As indicated by the behavior of pursuers observed in Section III, each pursuer must exhibit two distinct behaviors. When there is no nearby colliding pursuer, the $i$th pursuer’s objective is to reach the desired proximity from the target to enclose it. However, if there is a nearest colliding pursuer, the $i$th pursuer needs to maneuver along a loiter circle with a radius larger than the one corresponding to the nearest colliding pursuer to prevent any potential collisions. These factors enable us to break down the complex multiagent target enclosing problem into more manageable two-body (one pursuer and the target) or three-body (two pursuers and the target) scenarios at any given point in time. Consequently, in this subsection, we present the design of a potential function that will serve as the foundation for developing the target enclosing guidance law.
Let’s define the range error for the $i$th pursuer as,
${e}_{i}=r_{iT}-r_{d},$ (22)
where $r_{d}$ is the desired fixed proximity from the target such that $\dot{r}_{d}=0$. This error variable plays a crucial role in the design of the potential function to ensure inter-pursuer safety. To analyze the error dynamics, we differentiate $e_{i}$ with respect to time and use (1a) to obtain the range error rate for the $i$th pursuer as,
$\dot{e}_{i}=\dot{r}_{iT}=-v\cos\sigma_{i}$ (23)
since $\dot{r}_{d}=0$. From the above equation, it is evident that $\dot{e}_{i}<0,$ if $\sigma_{i}\in[0,\pi/2)$, which indicates the $i$th pursuer’s motion toward the target, whereas $\dot{e}_{i}>0,$ represents the opposite case because $\sigma_{i}\in(\pi/2,\pi]$. The $i$th pursuer moves on a circle of fixed radius from the target when $\dot{e}_{i}=0$ when $\sigma_{i}=\pi/2$. Therefore, nullifying the range error given in (22), will ensure that the $i$th pursuer converges to the desired proximity around the target, but additional safety considerations are needed to prevent the inter-pursuer collision.
We propose a potential function for the $i$th pursuer that incorporates both the required behaviors, which is defined as
$U_{i}(e_{i},e_{j})=\lambda_{i}\frac{e_{i}^{2}}{2}+\eta_{i}\delta_{i}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}$ (24)
where $\lambda_{i}$, $\eta_{i}$, and $\Delta_{i}$ are positive scaling constants in the potential function. The subscript $j$ corresponds to $P_{j}\in\mathcal{C}_{i}$, which is the nearest colliding pursuer corresponding to the $i$th pursuer. The variable $\delta_{i}$ is a conditional one such that
$\delta_{i}=\begin{cases}1;\,&\text{if}\,\mathcal{C}_{i}\neq\emptyset~{}(\mathcal{C}_{i}~{}\text{is not empty}),\\ 0;\;&\text{otherwise}.\end{cases}$ (25)
The first term in (24) is the attractive component of the potential function, which is active at all times. This term has a unique minimum at $e_{i}=0$ and steers the $i$th pursuer to the desired proximity around the target. The second term in (24) is the repulsive term enabling the $i$th pursuer to avoid collisions with its nearest colliding pursuer. It is important to note that the latter term is conditional and is only activated based on the existence of the nearest colliding pursuer for $P_{i}$, as per the condition given in (25). The first term in (24) is a quadratic term dependent on $e_{i}$, whereas the second term (which introduces a repulsive behavior) is a Gaussian function that attains maximum values at $e_{i}=e_{j}$ and decreases as $d_{ij}=e_{i}-e_{j}$ increases, with this term losing its dominance when $d_{ij}\geq\Delta_{i}$. If two agents are closer to each other, it is reasonable to assume $d_{ij}<e_{i}$. Therefore, the first term in (24) will dominate the second term for any selected value of $\lambda_{i}$, $\eta_{i}$ and $\Delta_{i}$ if $d_{ij}\gg 0$. However, for small values of $d_{ij}$, one can select the values of these parameters to ensure that the second term dominates in the region where $d_{ij}=e_{i}-e_{j}\in(0,\Delta_{i})$. In such a scenario, in the region near the line $e_{i}-e_{j}=0$, the repulsive term dominates the attractive component that will result in $P_{i}$ moving away from $P_{j}\in\mathcal{C}_{i}$. Consequently, it follows that there will be a shift in the minimum of the potential function at $e_{i}=0$ (for only attractive potential) to $e_{i}>e_{j}$ (for combined potential), which allows pursuers to avoid collisions with each other.
To analyze the combined potential function, we compute partial derivatives of (24) with respect to $e_{i}$ and $e_{j}$ to obtain the gradient of the potential function as,
$\frac{\partial U_{i}}{\partial e_{i}}$ $=\lambda_{i}e_{i}-\frac{\eta_{i}\delta_{i}}{\Delta^{2}}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}(e_{i}-e_{j}),$ (26a)
$\frac{\partial U_{i}}{\partial e_{j}}$ $=\frac{\eta_{i}\delta_{i}}{\Delta^{2}}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}(e_{i}-e_{j}).$ (26b)
To obtain the region of minimum potential, we equate (26a) to zero to obtain, that is, $\lambda_{i}e_{i}-\frac{\eta_{i}\delta_{i}}{\Delta^{2}}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}(e_{i}-e_{j})=0.$ This above expression is highly nonlinear, and the analytical solution of $e_{i}$ is difficult to obtain in closed form. The trivial solution is $e_{i}=0$ and $e_{j}=0$. However, such a situation never arises because when $\delta_{i}=1$, $e_{i}>e_{j}$. Further, from the above equation, we can infer that the solution to $\dfrac{\partial U_{i}}{\partial e_{i}}=0$ and $\dfrac{\partial U_{i}}{\partial e_{j}}=0$, can be given by $e_{i}-e_{j}=\epsilon_{i}$, where $\epsilon_{i}$ denotes a positive value. To verify this, we plot the potential function and the negative gradients for the chosen (typical) values of $\lambda_{i}=0.2,\eta_{i}=100000$ and $\Delta_{i}=500$, as depicted in Figure 9. In Figure 9, $e_{i}-e_{j}>0$ is the relevant area whenever $\delta_{i}=1$. The combined potential is depicted in Figure 8(a), where it is observed that the minimum potential region (depicted by cyan colored region) is situated at $e_{i}-e_{j}=\epsilon_{i}$. Further, in Figure 8(b), the negative gradient lines converge to the line $e_{i}-e_{j}=\epsilon_{i}$. This depicts the existence of a minimum potential point at $e_{i}=e_{j}+\epsilon_{i}$, whenever the repulsive component of the potential is activated.
Remark 5.
It is crucial to choose the values of the control parameters $\lambda_{i}$, $\eta_{i}$, and $\Delta_{i}$ judiciously to guarantee that $\epsilon_{i}>0$. This ensures that the minimum of the combined potential function in (24) occurs at $e_{i}>e_{j}$, leading to $\dot{e}_{i}>0$ for all $e_{i}-e_{j}\in(0,\epsilon_{i})$ and $\dot{e}_{i}<0$ for all $e_{i}-e_{j}\in(\epsilon_{i},\infty)$. Based on (26a), the condition, $\dfrac{\delta_{i}}{\Delta_{i}^{2}\lambda_{i}}>\sup_{t\geq 0}\frac{e_{i}}{e_{i}-e_{j}}e^{\left(\frac{\left(e_{i}-e_{J}\right)^{2}}{2\Delta_{i}^{2}}\right)}$, will help select the control parameters to ensure $\epsilon_{i}>0$.
(a) Potential function $\mathcal{U}_{i}$.
(b) Gradient lines.
Figure 9: Potential function and gradient lines for the $P_{i}$ in presence of a nearest colliding pursuer $P_{j}$, with $\lambda_{i}=0.2,\eta_{i}=100000$ and $\Delta_{i}=500$.
III-C Development of the Target Enclosing Guidance Law
In this section, we design the lateral acceleration (control input) for the $i$th pursuer to steer it on a collision-free path to the desired proximity around the target. To achieve this, we leverage the sliding mode control avowed for its remarkable properties of precision, robustness, and ease of designing, e.g., see [28, 29, 6, 21, 22, 23] .
Lemma 2.
The dynamics of the $i$th pursuer’s range error has a relative degree of two with respect to its lateral acceleration.
Proof.
Differentiating $\dot{e}_{i}$ with respect to time and using (1b), we obtain the dynamics of range error rate as,
$\ddot{e}_{i}$ $=\ddot{r}_{iT}=\dfrac{v^{2}\sin^{2}{\sigma_{i}}}{r_{iT}}+a_{i}\sin\sigma_{i},$ (27)
thereby completing the proof. ∎
Inferences from Lemma 2 allow us to construct a suitable sliding manifold for the $i$th pursuer. To this end, consider the sliding manifold,
$S_{i}={\dot{e}}_{i}+\lambda_{i}{e}_{i}-\frac{\eta_{i}\delta_{i}}{\Delta_{i}^{2}}\exp\left\{{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta_{i}^{2}}}\right\}({e}_{i}-{e}_{j}),$ (28)
on which the $i$th pursuer’s range error rate $\dot{e}_{i}$ can be driven to the negative gradient $-\nabla_{e_{i}}\mathcal{U}_{ij}$. Note from the above equation that when $\delta_{i}=0$ $S_{i}=\dot{e}_{i}+\lambda_{i}e_{i}$, which is analogous to a second-order sliding manifold with asymptotic error convergence. Further, the last term in (28) represents the collision avoidance term that becomes active when $\delta_{i}=1$. We now propose the lateral acceleration (the only control input) for the $i$th pursuer as,
$a_{i}=$ $\dfrac{1}{\sin\sigma_{i}}\Bigg{[}-K_{i}\mathrm{sign}({S_{i}})-\frac{v^{2}\sin^{2}{\sigma_{i}}}{r_{iT}}+\lambda_{i}v\cos\sigma_{i}$
$+\frac{\delta_{i}\eta_{i}ve^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\sigma^{2}}}}{\Delta_{i}^{2}}\left(\frac{({e}_{i}-{e}_{j})^{2}}{\Delta_{i}^{2}}-1\right)\cos\sigma_{i}\Bigg{]},$ (29)
$\textup{where},\,K_{i}>\frac{\eta_{i}v}{\Delta_{i}^{2}}\left(1-\frac{r_{s}^{2}}{\Delta_{i}^{2}}\right)$ (30)
provides a sufficient condition on the controller gain of the $i$th pursuer. The first term in the above equation drives the pursuers to their respective sliding manifold. The second term represents the centripetal acceleration about the target, while the third term represents the component of its velocity along its LOS with respect to the target. The fourth term in (III-C) enforces a collision avoidance behavior.
Theorem 2.
Consider the equations of relative motion between a single target and multiple pursuers (II-A), and the $i$th pursuer-target range error $e_{i}$. If the $i$th pursuer applies the lateral acceleration given in (III-C), then all the pursuers will self-organize on circle of radius $r_{d}$ around target without colliding with each other.
Proof.
Consider a Lyapunov Function Candidate $V=\sum_{i\in\mathcal{L}}S_{i}^{2}.$ Differentiating this Lyapunov function with respect to time and using (28), we obtain the dynamics of the sliding manifold as,
$\dot{V}$ $=\sum_{i\in\mathcal{L}}S_{i}\dot{S}_{i}=\sum_{i\in\mathcal{L}}S_{i}\left(\ddot{e}_{i}+\lambda_{i}\dot{e}_{i}+\frac{\partial^{2}U_{i}}{\partial^{2}{e}_{i}}\dot{e}_{i}+\frac{\partial^{2}U_{i}}{\partial^{2}{e}_{j}}\dot{e}_{j}\right)$ (31)
Substituting for $\ddot{e}_{i},\,\partial^{2}\mathcal{U}_{i}/\partial^{2}e_{i}$ and $\partial^{2}\mathcal{U}_{i}/\partial^{2}e_{j}$ using (26a), (26b), and (27) in the above expression we obtain,
$\dot{V}$ $=\sum_{i\in\mathcal{L}}S_{i}\Biggr{[}\frac{v^{2}\sin^{2}{\sigma_{i}}}{r_{iT}}+a_{i}\sin\sigma_{i}+\lambda\dot{e}_{i}$
$+\frac{\eta_{i}\delta_{i}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta^{2}}}}{\Delta_{i}^{2}}\left(\frac{({e}_{i}-{e}_{j})^{2}}{\Delta_{i}^{2}}-1\right)\left(\dot{e}_{i}-\dot{e}_{j}\right)\Biggr{]}.$ (32)
Substituting the proposed lateral acceleration for the $i$th pursuer given in (III-C), reduces the above expression to
$\dot{V}$ $=\sum_{j\in\mathcal{L}}S_{i}\Bigg{[}-K_{i}\mathrm{sign}(S_{i})-\frac{\eta_{i}\delta_{i}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta^{2}}}}{\Delta_{i}^{2}}\left(\frac{({e}_{i}-{e}_{j})^{2}}{\Delta_{i}^{2}}-1\right)\dot{e}_{j}\Bigg{]},$
$\leq\sum_{j\in\mathcal{L}}-|S_{i}|\Bigg{[}K_{i}-\frac{\eta_{i}}{\Delta_{i}^{2}}\bigg{|}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta^{2}}}\bigg{|}\,|\dot{e}_{j}|\,\Bigg{|}\frac{({e}_{i}-{e}_{j})^{2}}{\Delta^{2}}-1\Bigg{|}\Bigg{]},$
since $|\dot{e}_{j}|=v$ and $|e_{i}-e_{j}|<r_{s}$ (as the switching term can be activated only when the colliding vehicle is within the sensing range). The above equation can be further simplified as
$\dot{V}\leq\sum_{j\in\mathcal{L}}-|S_{i}|\Bigg{[}K_{i}-\frac{\eta_{i}v}{\Delta_{i}^{2}}\left(1-\frac{r_{s}^{2}}{\Delta_{i}^{2}}\right)\Bigg{]},$ (33)
which presents the sufficient condition as given in (30) to ensure, $\dot{V}<0,\,\forall\,S_{i}\in\mathbb{R}\setminus\{0\}$. This implies that $S_{i}$’s converge to zero within finite time given by $T^{*}=|S_{i}(0)|/K_{i}$, irrespective whether $\delta_{i}=0\,\textup{or}\,1$. It is worth mentioning that $S_{i}$ can suddenly deviate from zero due to the activation/deactivation of the repulsive potential component. However, based on (33), even when $S_{i}$ shifts suddenly to a non-zero value, $S_{i}$ will converge to a zero within a finite time to quickly enforce the sliding mode. Once $S_{i}=0$, the pursuers’ error rate propagates according to
${\dot{e}}_{i}=-\lambda_{i}{e}_{i}+\frac{\eta_{i}\delta_{i}}{\Delta_{i}^{2}}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta_{i}^{2}}}({e}_{i}-{e}_{j})=0.$ (34)
Therefore, once the sliding mode is enforced, $P_{i}$ may exhibit two distinct behaviors in relation to other vehicles based on the switching term, discussed as Modes 1 and 2 below.
Mode 1: When $\delta_{i}=0$, indicating the absence of any nearest colliding pursuer for $P_{i}$, then the dynamics described in (34) yields ${\dot{e}}_{i}=-\lambda_{i}{e}_{i}$. This leads to 1) ${\dot{e}}_{i}\leq 0$ when $e_{i}\geq 0$ and 2) ${\dot{e}}_{i}>0$ when $e_{i}<0$. The solution of the range error can obtained as $\;{e}_{i}(t)={e}_{i}(0)e^{-\lambda_{i}t}$. Consequently, the pursuer converges to the desired proximity, that is, $r_{iT}\rightarrow r_{d}$ if $\lambda_{i}$ is selected as a positive value.
Mode II: If ${\delta}_{i}=1$, this pertains to the scenario when there exists a nearest colliding neighbor $P_{j}$ for the pursuer $P_{i}$ such that $d_{ij}>0$ and $\psi_{ij}<0$ at the start of the engagement. In this mode, (34) yields
${\dot{e}}_{i}=-\lambda_{i}{e}_{i}+\frac{\eta_{i}}{\Delta^{2}}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta^{2}}}({e}_{i}-{e}_{j})=0,$ (35)
as $\delta_{i}=1$. It can be inferred that the solution of $e_{i}(t)$ when $\delta_{i}=1$ may not be feasible to obtain analytically in each case. However, we can select the values of controller parameters $\lambda_{i}$, $\eta_{i}$ and $\Delta_{i}$ to ensure $\dot{e_{i}}>0$ when $e_{i}-e_{j}<\epsilon_{i}$ and $\dot{e_{i}}<0$ when $e_{i}-e_{j}>\epsilon_{i}$, where $\epsilon_{i}>0$ denotes a positive value. This is similar to the behavior of the pursuer under the proposed potential function as discussed earlier in Section III-B. This implies $P_{i}$ will converge to a loiter circle of radius greater than that of $P_{j}$, that is, $e_{i}\to e_{j}+\epsilon_{i}\,\,\textup{or}\,\,r_{iT}\to r_{jT}+\epsilon_{i}.$ It readily follows from the above relation that $d_{ij}>0$, which represents the safety condition as in (8). The above condition ensures $d_{ij}>0,\,\,\forall\,t\geq 0$, while $\psi_{ij}<0$ and $\dot{\psi}_{ij}>0$ in Region IV of the $i$th pursuer, as obtained in (III-A) of Lemma 1. This will result in $\psi_{ij}$ starting from a negative value to become zero. At this moment, $P_{j}$ will no longer be the nearest colliding neighbor for $P_{i}$. Further, if no other vehicle becomes the new nearest colliding neighbor, $P_{i}$ will undergo the behavior in Mode I to reach the desired proximity, settling behind $P_{j}$ and exhibiting the self-organizing behavior. Otherwise, $P_{i}$ will undergo the behavior in Mode II to avoid collision with the new nearest colliding neighbor and settle behind it. Consequently, every pursuer starting away from the target will alternate between Mode I and Mode II behaviors, resulting in every pursuer to converge to circle of desired proximity without colliding with other pursuers and organising autonomously on enclosing shape. This concludes the proof. ∎
The proposed guidance law (III-C) is nonsingular even if the term $\sigma_{i}$ appears in its denominator, that is, the expression in (III-C) never becomes unbounded if $\sigma_{i}\rightarrow 0\,\textup{or}\,\,\pm\pi$. In the steady state, $\dot{e}_{i}=-v\cos\sigma_{i}=0$ results in $\sigma_{i}\to\pm\pi/2$ as the equilibrium points. Therefore, $\sigma_{i}=0\,\textup{or}\,\pi$ are not the equilibrium points for the pursuers utilizing the proposed guidance law. However, $\sigma_{i}=0\,\,\textup{or}\,\,\pi$ may occasionally occur in the transient phase during which $a_{i}$ will saturate momentarily to drive the system away from such configurations. However, during implementation, control inputs are constrained to remain within finite bounds since the control input values practically cannot be infinite. Therefore, whenever the demand for $a_{i}$ grows beyond the allowable limits, the control input will saturate and steer the pursuer away from $\sigma_{i}=0\,\,\textup{or}\,\,\pi$.
Remark 6.
The control input designed in Theorem 2 is solely based on relative measurements, making it lucrative even in scenarios when global information is not available. The proposed guidance law for the $i$th pursuer (III-C) is independent of the other pursuers’ motion, thereby exhibiting robustness to their movement.
Remark 7.
Ensuring the expression of $K_{i}$ as in (30) is greater than zero provides an additional condition on the selection of the control parameters, that is, $\Delta_{i}>r_{s}$. This simply implies that the region of the dominance of the repulsive term denoted by $\Delta_{i}$ is greater than the $i$th pursuer’s sensing radius, resulting in the repulsive field dominating within the sensing radius. This will enable the pursuers to move out of the sensing range of their respective nearest colliding pursuers to avoid colliding with them, such that at steady state, none of the pursuers are situated in each other’s sensing radius.
IV Simulation Results
In this section, we demonstrate the efficacy of the proposed control law in autonomously organizing the pursuers to enclose a stationary target. In the following results, the target is stationary at the position $[x_{T}\;y_{T}]=[0,0]^{\top}$ m. All pursuers start at a different distances from the target and move at a constant speed of $40$ m/s and. The initial position of pursuers are denoted by square markers in the subsequent trajectory plots. The desired proximity from the target is set as $r_{d}=100$ m for all the simulation results. The controller gains and other parameters are selected as follows: $K_{i}=10$, $\lambda_{i}=0.9$, $\eta_{i}=70000$, $\Delta=100$, and $r_{s}=50$ m.
(a) Trajectories.
(b) Error variables and sliding manifold.
(c) Lateral accelerations.
(d) Inter-pursuer distances and angular spacings.
Figure 10: Collision-free self-organising target enclosing with 6 pursuers.
The first set of results illustrates the scenario in which six pursuers start away from the target, as depicted in Figure 10. The initial arrangement of the pursuers around the target in a clockwise direction is given as $P_{2}-P_{1}-P_{3}-P_{4}-P_{5}-P_{6}$, as shown by square markers in the figure. It is observed that the final arrangement of the pursuers on the desired proximity (as depicted by circular markers) is given as, $P_{2}-P_{3}-P_{1}-P_{4}-P_{6}-P_{5}$. This indicates that only $P_{3}$ and $P_{6}$ encounter a nearest colliding pursuer and reorganize to avoid collision. Meanwhile, the $P_{1},P_{2},P_{4},andP_{5}$ move directly to the desired proximity without encountering the nearest colliding pursuer. The profiles of the error variables and sliding manifold are depicted Figure 9(b), where the sliding manifold converges to zero within a finite time. The profiles of $S_{i}$ for only $P_{3}$ and $P_{6}$ depict sudden deviations at different instances of time, which represent the instance of activation/deactivation of the collision avoidance component of the control input. Note that whenever collision avoidance is activated, the range error diverges from zero momentarily to ensure safety.
Figure 9(c) depicts the profiles of lateral accelerations of the pursuers converging to the same constant value at a steady state, as all the pursuers eventually converge to the same loiter circle of radius of desired proximity. The lateral acceleration profiles of $P_{3}$ and $P_{6}$ vary from the other pursuers due to the activation of the collision avoidance component. Finally, the profiles of inter-pursuer variables are shown in Figure 9(d), where the profiles of $\psi_{ij}$ change sign crossing zero for only $P_{3}$ and $P_{6}$. This shows that these pursuers have autonomously reorganized from their initial configurations, where the profiles $r_{ij}^{\min}$ depict that the distance between the two closest pursuers never reaches zero. Further, at steady state, all the pursuers have $r_{ij}^{\min}>r_{s}$, which implies that all the pursuers only converge to the desired proximity when the sensing range of no two pursuers intersect. Therefore, all the pursuers have only the attractive component of the control input activated at the steady state.
Figure 11 illustrates the second set of results where initially, four pursuers ($P_{1}$ to $P_{4}$) are already moving on the circle of desired proximity when two other pursuers ($P_{5}$ and $P_{6}$) start away from the desired proximity. From the Figures 10(a), 10(b) and 10(c), it is evident that $P_{5}$ reaches the desired proximity without encountering a nearest colliding pursuer and autonomously organizing on the free space on the enclosing shape. While $P_{5}$ itself becomes the nearest colliding pursuer for $P_{6}$, resulting in $P_{6}$ taking a larger route to converge to the desired proximity behind $P_{6}$. This is verified by Figure 10(d), where only the profiles related to $P_{6}$ exhibit abrupt variation compared to other pursuers, indicating to activation of the repulsive component of the control input. Therefore, $P_{5}$ and $P_{6}$ are able to autonomously join the enclosing formation without colliding with other pursuers.
Furthermore, we offer video demonstrations of our simulation results, accessible at this link. The videos titled “results_1.mp4” and “results_2.mp4” showcase the simulation outcomes corresponding to the scenarios presented in the paper. Additionally,“results_3.mp4” and “results_4.mp4” illustrate the self-organizing target enclosing in large swarms with 10 and 20 pursuers, respectively.
(a) Trajectories ($t=0\,sec$).
(b) Trajectories ($t=15\,sec$).
(c) Trajectories ($t=35\,sec$).
(d) Lateral accelerations.
Figure 11: Two pursuers joining the enclosing formation.
V Conclusions
In this paper, we proposed guidance laws for multiple pursuers to safely and autonomously organize at the desired proximity to encircle a stationary target, using only the relative information. In our approach, the agents are free to converge to any point on the enclosing shape as long as they maintain a safe distance between them, which may result in the pursuers not requiring a pre-determined formation structure and autonomously organizing on the enclosing shape. Our guidance strategy ensures that each pursuer selects at most one other pursuer based on certain collision conditions. We show that preventing collision with the selected other pursuer is sufficient to avoid collisions with all the other pursuers. To tailor the desired switching behavior required, we proposed a potential function such that the attractive component of the potential is activated all the time. Meanwhile, the repulsive component is activated whenever the pursuer encounters a nearest colliding pursuer. Further, we utilized sliding mode control to develop the reaching law, ensuring finite time convergence of the sliding manifold, followed by asymptotic convergence of the error to zero. The resulting guidance law enables pursuers to move towards the desired proximity along collision-free paths, offering increased maneuvering flexibility, requiring minimal information, and demonstrating robustness to the motion of other pursuers. Validation of the guidance laws with different numbers of pursuers attests to the merits of the proposed guidance algorithm. Potential future extensions of this work may involve incorporating a moving target, exploring three-dimensional spaces, and integrating obstacle avoidance techniques.